THEORETICAL AND EXPERIMENTAL STUDY OF THE INFRARED SPECTRA OF CO_{2}-(para-H_{2})_{2} TRIMERS

\underline{HUI LI}, Department of Chemistry, University of Waterloo, Waterloo, Ontario N2L 3G1, Canada; A. R. W. McKELLAR, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, Ontario K1A OR6, Canada; ROBERT J. LE ROY, and PIERRE-NICHOLAS ROY, Department of Chemistry, University of Waterloo, Waterloo, Ontario N2L 3G1, Canada.

The infrared spectra of several isotopologues of the CO_{2}-(para-H_{2})_{2} trimer have been predicted by applying exact basis-set calculations to a global potential energy surface which is defined as a sum of the accurately known two-body para-H_{2}–CO_{2} and para-H_{2}–para-H_{2} potentials.^{a} These results are compared with new spectroscopic measurements, for which sixteen transitions have been assigned. A reduced-dimension treatment of the para-H_{2} rotation has been employed by performing an exact hindered-rotor average.^{b} Three-body effects and the quality of the potential are discussed. We present a new technique for representing the three-dimensional para-H_{2} density in the body-fixed frame. It shown the two para-H_{2} molecules are localized much more closely together than is the case for the two He atoms in the analogous CO_{2}–(He)_{2} system.

a H. Li, P.-N. Roy and R.J. Le Roy, J. Chem. Phys. (2010, submitted); K. Patkowski, W. Cencek, P. Jankowski, K. Szalewicz, J.B. Mehl, G. Garberoglio and A.H. Harvey, J. Chem. Phys. 129, 094304 (2010).

bH. Li, P.-N. Roy and R.J. Le Roy, J. Chem. Phys. (2010 submitted).